OpenStudy (anonymous):

z^2+iz-6z+10 What are the solutions for z? Note i is the imaginary number...

OpenStudy (anonymous):

What the hell?It is not an equation at all.But if you mean z^2+iz-6z+10=0,just make it to be z^2-6z+10=-iz.Then take the square of it:z^4+36z+100-12z^3+20z^2-120z=-z^2...Then solve this four times equation...good luck

jhonyy9 (jhonyy9):

so one different opinion yes this is right you can solve it if you write there in style of equation mated equal with zero z^2 +iz -6z +10 =0 and now you can factorizing out from middle terms the ,,z" and will get an quadratic equation in ,,z" what you need to solve it using the discriminant formula and continue it how you have learned in your school with z_1,2 = ... ok ? wait your reply soon with results good luck hope so much that is understandably bye